A recent poll pertaining to educational reforms involved 634 men and 820 women. The margin of error is 3.6% for the male sample and 3.2% for the female sample. Find the level of confidence for the male portion of the poll and for the female portion of the poll. HINT: See Example 7. (Round your answers to one decimal place.)
a) male %
b) female %

Answer by stanbon(75887)

A recent poll pertaining to educational reforms involved 634 men and 820 women. The margin of error is 3.6% for the male sample and 3.2% for the female sample. Find the level of confidence for the male portion of the poll and for the female portion of the poll. HINT: See Example 7. (Round your answers to one decimal place.)
a) male %
ME = z*sqrt[pq/n]
-----
0.036 = z*sqrt[(1/2)(1/2)/634]
0.036 = z(1/2)(1/25.18]
z/2 = 25.18*0.036
z = 1.81
---
normalcdf(1.81,100) = 0.035
level of confidence = 2*3.5 = 7%
------------------------------------

b) female %
z*(1/2)[1/sqrt(820)] = 0.032
----
z*(1/2)[1/28.64] = 0.032
----
z = 0.032*2*28.64 = 1.84
----
normalcdf(1.84,100) = 0.033
level of confidence = 2*3.3% = 6.6%
==========
Cheers,
Stan H.
---------------